A Pole 24 Feet Tall Has A Shadow Of Eight Feet Long. A Near By Pole Is 72feet Tall How Long Is The Shadow?
Answer by checkley77(12569)

24/8=72/x cross multiply
24x=8*72
24x=576
x=576/24
x=24 ft. is the length of the second shadow.
Proof:
24/8=72/24
3=3